OpenStudy (anonymous):

MEDALS WILL BE GIVEN!!! The cubs have won 44 baseball games and have lost 65 games.They have 53 games remaining at least how many of the remaining 53 games must the cubs win to have a winning season ? (a winning season means they win more than 50% of their games)

OpenStudy (anonymous):

@ASAAD123 @Emah @emorainbow @Edward1992 @Frostbite @itsmylife @imakoolkid @josephniel14 @Math4Life @ObeyChris{~} @ParthKohli @saifoo.khan

OpenStudy (theviper):

Total Matches=44+65+53=162 Its 50% is 81 Therefore the club should win 81 matches & it has already win 44 matches Therfore club should win 81-44=? in the remaining 53 matches :)

OpenStudy (theviper):

not mentioned me :(

OpenStudy (anonymous):

CHOICES a: at least 54.5 games b: at least 38 games c: at least 53 games d: at least 26.5 games

OpenStudy (anonymous):

@TheViper

OpenStudy (anonymous):

and c

OpenStudy (theviper):

b is the right one

OpenStudy (theviper):

:DD

OpenStudy (theviper):

But now you got ur answer :)

OpenStudy (anonymous):

what do you mean you said that they need to win the remaining 53

OpenStudy (theviper):

I did not said that. I said that the should in 81-44=37+1=38(because they need to win more than 50%)

OpenStudy (anonymous):

oh ok sorry

OpenStudy (theviper):

np :))